OpenStudy (anonymous):

Select all of the following that are ordered pairs of the given function. f(x) = 3 - 2x (-2, -1) (-1, 5) (0, 3) (1, 0) (2, -1)

OpenStudy (phi):

you should replace x with the first number of the pair, do the arithmetic and see if the answer matches the 2nd number in the pair. if it does, that pair "works"

OpenStudy (ivettef365):

just substitute the points to see which one is true, example you have point (-2,-1) equation F(x) = 3 - 2x -1 = 3 - 2(-2) -1 = 3 + 4 -1 = 7 NOT TRUE